OpenStudy (anonymous):

a challenge question for all if 1035 + summation( tan( j pi/180) ) from j =0 to 45 = 2^x then x= 1) 20 2)21 3)22 4)23

zepdrix (zepdrix):

I don't know the answer, but I thought I'd format this for you so it's a little easier to read. Hopefully I read it correctly.\[\huge 1035+ \sum_{\color{royalblue}{j}=0}^{45}\; \tan\left(\color{royalblue}{j}\frac{\pi}{180}\right) \quad = \quad 2^x\] \[\huge x= \;\color{green}{?}\]

OpenStudy (compassionate):

x = 22

OpenStudy (anonymous):

@Compassionate nopes..though you were close on the logic i guess.. @zepdrix yep,,thats the right interpretation..

OpenStudy (anonymous):

how can it be a power of 2 with ur options :) because\[\large \sum_{\color{royalblue}{j}=0}^{45}\; \tan\left(\color{royalblue}{j}\frac{\pi}{180}\right) \quad <45 \ \tan \frac{\pi}{4}=45 \]if im not wrong

OpenStudy (anonymous):

please whosover sees it atleast bookmark it, so that if someone tells the solution or finally i tell the solution, you get to know that the solution really exists and is a genius one! :) note that i couldnt solve it either, but one of my professors did..

OpenStudy (anonymous):

not professors, teachers..

OpenStudy (zarkon):

\[N\left[1035+\sum _{j=0}^{45} \text{Tan}\left[j*\frac{\pi }{180}\right],20\right]\] Out[15]= 1055.3586583514306367

OpenStudy (shubhamsrg):

*

OpenStudy (anonymous):

@mukushla yep that inequality is correct but how do you come to your conclusion ? as according to me,,the question is perfectly fine..

OpenStudy (anonymous):

@mukushla has proved your question wrong with his Extra ordinary thinking.. ha ha ha...

OpenStudy (anonymous):

well i'd be glad if i am proved wrong..am waiting for him to reason this out..

OpenStudy (anonymous):

He comes rarely online now a days ..

OpenStudy (queelius):

Uh, I got x=10.0435176596052.

OpenStudy (anonymous):

So, it means the term : \(\large \sum_{\color{royalblue}{j}=0}^{45}\; \tan\left(\color{royalblue}{j}\frac{\pi}{180}\right)\) can have one of the four values according to the choices given : 1047541 Or 2096117 Or 4193269 Or 8387573..

OpenStudy (anonymous):

Something looks wrong to me..

OpenStudy (anonymous):

Do you have given right choices there @Aryang

OpenStudy (anonymous):

now that all of you are saying it,,even am smelling something fishy here..i'll confirm and get back here asap..

OpenStudy (anonymous):

It is better now to confirm it..

OpenStudy (anonymous):

is it an IIT question ?

OpenStudy (shubhamsrg):

well one thing i can easily make out is 1035 = summ( j) from j=0 to 45 so LHS on whole can be written as summation (j + tan(j pi/180) ) from j=0 to 45 my personal query is that, are you sure its summation there? if thats the product symbol there,,i can easily come to ans.otherwise i doubt,.,

OpenStudy (anonymous):

I got\[x \approx 10.04351766\]

OpenStudy (callisto):

*

OpenStudy (anonymous):

as @mukushla said  and there is no number which can be expressed as 2^x where x is a integer between 1035 and 1080